OpenStudy (anonymous):

Need help with Chemical Quantities. Calculate how many moles of products would be produced if .5 mole of the first reactant where to react completely? CO2+4 H2 -> CH4+2 H20

OpenStudy (aaronq):

Use the stoichiometric coefficients to find moles produced. Set up a ratio using the species of interest, like so: e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients , \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) From here you can isolate what you need. For example: if you have 2 moles of B, how many moles of C can you produce? solve algebraically: \(\dfrac{2}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\rightarrow n_C=\dfrac{2*\color{green}{c}}{\color{blue}{b}}\)